# Unsupervised and Source-Free Ranking of Biomedical Segmentation Models

### v2

@@ 8.3 Light Microscopy (LM), Cells
8.4 Cone-Beam Computed Tomography (CBCT), ToothFairy2
Multiclass Semantic Segmentation. For multiclass semantic segmentation experiments, we use ToothFairy2 from a MICCAI2024 challenge [9, 10]. It comprises 530 volumetric scans with voxel-wise annotations of maxillofacial anatomy, of which 480 are publicly available for training, while 50, acquired by a different institution, are held out for evaluation via the grand-challenge platform. The dataset provides annotations for 42 classes, including jawbones, left and right inferior alveolar canals, maxillary sinuses, pharynx, upper and lower teeth, bridges, crowns, and implants. For evaluation, we obtained access to the private test set, collected by a different centre on different acquisition machines.
9 Models
We trained a range of model architectures to build a diverse model set for transfer comparison. All of the models trained by us are trained on the training set of a single ‘source’ dataset. We additionally investigated several publicly available models taken from the Toothfairy2 challenge [9], the Bioimage Model Zoo [49] and large generalist models taken from popular repositories [51, 3]. Information listing the full set of models considered can be found in Sec. 4.2 in the main paper. In the following section we detail the model training setup.
@@ 9 Models
10 Model Perturbation
@@
Feature Space Perturbations: We relied upon test-time application of spatial DropOut [66] to provide feature space perturbation. We tested several approaches for applying DropOut to our models. Firstly, we tested applying Spatial DropOut only to the bottleneck layer of U-Net style architectures. Our experiments for bottleneck only DropOut are shown on the nuclei datasets both for semantic and instance segmentation (Tabs. 11, 18 and 32). Secondly, in addition to the bottle neck layer we also investigated adding DropOut to all layers that connect the encoder to the decoder, namely skip connections in U-Net style networks. Our experiments for bottleneck + skip connection DropOut are shown on the Toothfairy2 dataset for multiclass semantic segmentation (Tab. 21). Lastly, we investigate applying DropOut uniformly to every layer of the network, this represents the most general and architecture agnostic approach considered. Our experiments for all layer DropOut are shown on the mitochondria semantic segmentation datasets along with the cell instance segmentation datasets (Tabs. 7, 8 and 27).
@@
We found that all the considered feature space perturbation approaches were able to perform effectively, as shown in the tables in the following sections. However, it is worth noting that the first approach of applying DropOut to the bottleneck layer only can be sensitive to residual connections. The second approach of applying DropOut to both the bottleneck layer and skip connections works very well on the model sets tested, but does assume that the models have skip connections in the architecture, otherwise it reduces to the first case (bottleneck only) and could suffer the same sensitivity to residual connections. The last approach of applying DropOut to all layers is the most architecture agnostic and resulted in very strong correlation between the CTE metric and performance score.
11 Semantic Segmentation
@@
In the following section we investigate ranking of semantic segmentation models, each of the tables Tabs. 5, 7, 9, 11, 6, 8, 10 and 18 shows the correlation scores achieved by the CTE metrics for sets of models applied directly to a single target dataset, across a range of perturbation strengths. In the main paper, due to space constraints, we show the correlation scores for a single perturbation strength averaged over the set of target datasets.
@@
We investigate the dependency of CTE on the strength of perturbation applied for both input and feature space perturbations. The tables show that CTE’s performance is stable across a wide range of perturbations.
11.1 Mitochondria Perturbation Sweep
@@
For mitochondria semantic segmentation we investigated applying additive Gaussian noise as an input perturbation between the strengths . We show the results for both CTE-EI (Tab. 5) and CTE-NHD (Tab. 6).
@@
For feature space perturbation we investigated applying spatial DropOut equally to all layers of a network, where the strength of the perturbation is controlled by the proportion of feature maps dropped at each layer . We investigated DropOut proportions in the range . We show results for for both CTE-EI (Tab. 7) and CTE-NHD (Tab. 8). The tables show that CTE’s performance is stable across a wide range of perturbations.
@@
11.2 Nuclei Perturbation Sweep
For nuclei semantic segmentation we investigated applying a range of input augmentations; additive Gaussian noise, Gamma correction, Brightness adjustment and Contrast adjustment (as defined in Sec. 10). We investigated additive Gaussian noise with strengths in the range , Gamma correction with strengths in the range , where equals no adjustment, Brightness adjustment with strengths in the range and Contrast adjustment with strengths in the range , where equals no adjustment. We show the results for both CTE-EI (Tabs. 9, 14, 12 and 13) and CTE-NHD (Tabs. 10, 17, 15 and 16).
For feature space perturbation we investigated applying spatial DropOut only to the bottle neck layer of networks, where the strength of the perturbation is controlled by the proportion of feature maps dropped at each layer . We investigated DropOut proportions in the range . We show results for for both CTE-EI (Tab. 11) and CTE-NHD (Tab. 18).
| Metric | BBBC039 | DSB2018 | Hoechst | S_BIAD895 | |||||
@@ 11.2 Nuclei Perturbation Sweep
| Metric | BBBC039 | DSB2018 | Hoechst | S_BIAD895 | |||||
@@
| Metric | BBBC039 | DSB2018 | Hoechst | S_BIAD895 | |||||
@@
| CTE-EI DropOut ( 0.05) | K | 0.71 | (*) | 0.43 | (0.24) | 0.52 | (0.18) | 0.60 | (0.13) |
@@
| Metric | BBBC039 | DSB2018 | Hoechst | S_BIAD895 | |||||
@@
| CTE-EI Ctr ( 0.8 - 0.9) | K | 0.81 | (*) | 0.33 | (0.35) | 0.52 | (0.16) | 0.47 | (0.32) |
| S | 0.93 | (*) | 0.50 | (0.26) | 0.68 | (0.12) | 0.71 | (0.11) | |
| P | 0.95 | (**) | 0.91 | (**) | 0.94 | (**) | 0.98 | (**) | |
| CTE-EI Ctr ( 0.9 - 0.95) | K | 0.71 | (*) | 0.33 | (0.36) | 0.52 | (0.15) | 0.47 | (0.26) |
| S | 0.86 | (*) | 0.50 | (0.27) | 0.68 | (0.12) | 0.71 | (0.12) | |
| P | 0.98 | (**) | 0.95 | (**) | 0.95 | (**) | 0.98 | (**) | |
| CTE-EI Ctr ( 1.05 - 1.1) | K | 0.62 | (0.09) | 0.43 | (0.28) | 0.62 | (0.08) | 0.47 | (0.27) |
| S | 0.71 | (0.10) | 0.64 | (0.13) | 0.75 | (0.09) | 0.71 | (0.13) | |
| P | 0.98 | (**) | 0.97 | (**) | 0.94 | (**) | 0.98 | (**) | |
| CTE-EI Ctr ( 1.1 - 1.2) | K | 0.62 | (0.07) | 0.43 | (0.26) | 0.62 | (0.09) | 0.47 | (0.29) |
| S | 0.71 | (0.11) | 0.64 | (0.15) | 0.75 | (0.05) | 0.71 | (0.15) | |
| P | 0.99 | (**) | 0.97 | (**) | 0.95 | (**) | 0.98 | (**) | |
| Metric | BBBC039 | DSB2018 | Hoechst | S_BIAD895 | |||||
| pval. | pval. | pval. | pval. | ||||||
| CTE-EI Gamma ( 0.8 - 0.9) | K | 0.71 | (*) | 0.52 | (0.12) | 0.81 | (**) | 0.60 | (0.12) |
| S | 0.86 | (*) | 0.68 | (0.13) | 0.89 | (*) | 0.77 | (0.09) | |
| P | 0.98 | (**) | 0.96 | (**) | 0.98 | (**) | 0.97 | (**) | |
| CTE-EI Gamma ( 0.9 - 0.95) | K | 0.71 | (*) | 0.52 | (0.14) | 0.71 | (*) | 0.47 | (0.26) |
| S | 0.86 | (*) | 0.68 | (0.13) | 0.82 | (*) | 0.71 | (0.16) | |
| P | 0.99 | (**) | 0.96 | (**) | 0.98 | (**) | 0.96 | (**) | |
| CTE-EI Gamma ( 1.05 - 1.1) | K | 0.71 | (*) | 0.52 | (0.13) | 0.62 | (0.06) | 0.47 | (0.26) |
| S | 0.86 | (*) | 0.68 | (0.12) | 0.75 | (0.08) | 0.71 | (0.14) | |
| P | 0.98 | (**) | 0.94 | (**) | 0.97 | (**) | 0.97 | (**) | |
| CTE-EI Gamma ( 1.1 - 1.2) | K | 0.81 | (*) | 0.52 | (0.16) | 0.81 | (*) | 0.73 | (0.06) |
| S | 0.93 | (**) | 0.68 | (0.14) | 0.89 | (*) | 0.83 | (0.07) | |
| P | 0.96 | (**) | 0.94 | (**) | 0.98 | (**) | 0.96 | (**) | |
| Metric | BBBC039 | DSB2018 | Hoechst | S_BIAD895 | |||||
| pval. | pval. | pval. | pval. | ||||||
| CTE-EI Brt ( 0.0 - 0.05) | K | 0.71 | (*) | 0.43 | (0.25) | 0.52 | (0.16) | 0.60 | (0.12) |
| S | 0.86 | (*) | 0.64 | (0.15) | 0.71 | (0.08) | 0.77 | (0.08) | |
| P | 0.99 | (**) | 0.91 | (**) | 0.94 | (**) | 0.96 | (**) | |
| CTE-EI Brt ( 0.05 - 0.1) | K | 0.81 | (*) | 0.62 | (0.08) | 0.81 | (*) | 0.73 | (0.07) |
| S | 0.93 | (*) | 0.75 | (0.07) | 0.93 | (**) | 0.83 | (0.05) | |
| P | 0.99 | (**) | 0.99 | (**) | 0.94 | (**) | 0.99 | (**) | |
| CTE-EI Brt ( 0.1 - 0.2) | K | 0.71 | (*) | 0.62 | (0.06) | 0.81 | (*) | 0.73 | (0.06) |
| S | 0.86 | (*) | 0.75 | (0.07) | 0.93 | (*) | 0.83 | (0.07) | |
| P | 0.99 | (**) | 0.99 | (**) | 0.94 | (**) | 0.99 | (**) | |
| Metric | BBBC039 | DSB2018 | Hoechst | S_BIAD895 | |||||
| pval. | pval. | pval. | pval. | ||||||
| CTE-NHD Ctr ( 0.8 - 0.9) | K | 0.81 | (*) | 0.43 | (0.23) | 0.52 | (0.12) | 0.73 | (0.05) |
| S | 0.93 | (*) | 0.54 | (0.20) | 0.71 | (0.09) | 0.83 | (0.05) | |
| P | 0.94 | (**) | 0.91 | (**) | 0.92 | (**) | 0.99 | (**) | |
| CTE-NHD Ctr ( 0.9 - 0.95) | K | 0.81 | (*) | 0.43 | (0.25) | 0.81 | (*) | 0.60 | (0.12) |
| S | 0.93 | (*) | 0.54 | (0.24) | 0.93 | (*) | 0.77 | (0.09) | |
| P | 0.98 | (**) | 0.96 | (**) | 0.97 | (**) | 0.98 | (**) | |
| CTE-NHD Ctr ( 1.05 - 1.1) | K | 0.81 | (*) | 0.43 | (0.27) | 0.71 | (*) | 0.73 | (0.07) |
| S | 0.93 | (**) | 0.54 | (0.21) | 0.89 | (*) | 0.83 | (0.07) | |
| P | 0.99 | (**) | 0.97 | (**) | 0.92 | (**) | 0.99 | (**) | |
| CTE-NHD Ctr ( 1.1 - 1.2) | K | 0.81 | (*) | 0.43 | (0.24) | 0.71 | (0.05) | 0.60 | (0.14) |
| S | 0.93 | (*) | 0.64 | (0.12) | 0.89 | (*) | 0.77 | (0.08) | |
| P | 0.99 | (**) | 0.99 | (**) | 0.91 | (**) | 0.99 | (**) | |
| Metric | BBBC039 | DSB2018 | Hoechst | S_BIAD895 | |||||
| pval. | pval. | pval. | pval. | ||||||
| CTE-NHD Gamma ( 0.8 - 0.9) | K | 0.71 | (*) | 0.62 | (0.05) | 0.90 | (*) | 0.87 | (*) |
| S | 0.86 | (*) | 0.75 | (0.06) | 0.96 | (**) | 0.94 | (*) | |
| P | 0.97 | (**) | 0.91 | (**) | 0.99 | (**) | 0.95 | (**) | |
| CTE-NHD Gamma ( 0.9 - 0.95) | K | 0.62 | (0.07) | 0.52 | (0.14) | 0.81 | (*) | 0.87 | (*) |
| S | 0.75 | (0.06) | 0.68 | (0.13) | 0.93 | (*) | 0.94 | (*) | |
| P | 0.97 | (**) | 0.92 | (**) | 1.00 | (**) | 0.95 | (**) | |
| CTE-NHD Gamma ( 1.05 - 1.1) | K | 0.62 | (0.08) | 0.43 | (0.26) | 1.00 | (**) | 0.73 | (0.05) |
| S | 0.82 | (0.05) | 0.57 | (0.19) | 1.00 | (**) | 0.83 | (0.08) | |
| P | 0.93 | (**) | 0.88 | (*) | 1.00 | (**) | 0.83 | (*) | |
| CTE-NHD Gamma ( 1.1 - 1.2) | K | 0.71 | (*) | 0.62 | (0.06) | 0.81 | (*) | 1.00 | (**) |
| S | 0.79 | (*) | 0.75 | (0.06) | 0.93 | (*) | 1.00 | (*) | |
| P | 0.84 | (*) | 0.89 | (*) | 1.00 | (**) | 0.80 | (0.06) | |
| Metric | BBBC039 | DSB2018 | Hoechst | S_BIAD895 | |||||
| pval. | pval. | pval. | pval. | ||||||
| CTE-NHD Brt ( 0.0 - 0.05) | K | 0.52 | (0.12) | 0.71 | (*) | 0.71 | (*) | 0.47 | (0.25) |
| S | 0.71 | (0.08) | 0.79 | (0.05) | 0.86 | (*) | 0.54 | (0.29) | |
| P | 1.00 | (**) | 0.82 | (*) | 0.92 | (**) | 0.82 | (0.05) | |
| CTE-NHD Brt ( 0.05 - 0.1) | K | 0.71 | (*) | 0.62 | (0.07) | 0.71 | (*) | 0.73 | (0.05) |
| S | 0.86 | (*) | 0.71 | (0.08) | 0.86 | (*) | 0.83 | (0.06) | |
| P | 1.00 | (**) | 0.98 | (**) | 0.90 | (*) | 0.99 | (**) | |
| CTE-NHD Brt ( 0.1 - 0.2) | K | 0.71 | (*) | 0.71 | (*) | 0.90 | (*) | 0.73 | (0.07) |
| S | 0.86 | (*) | 0.82 | (*) | 0.96 | (**) | 0.83 | (0.07) | |
| P | 1.00 | (**) | 0.99 | (**) | 0.93 | (**) | 0.99 | (**) | |
| Metric | BBBC039 | DSB2018 | Hoechst | S_BIAD895 | |||||
| pval. | pval. | pval. | pval. | ||||||
| CTE-NHD DropOut ( 0.05) | K | 0.71 | (*) | 0.43 | (0.23) | 0.52 | (0.15) | 0.87 | (*) |

### v3

@@ 0.A.3 Light Microscopy (LM), Cells
0.A.4 Cone-Beam Computed Tomography (CBCT), ToothFairy2
Multiclass Semantic Segmentation. For multiclass semantic segmentation experiments, we use ToothFairy2 from a MICCAI2024 challenge [2024TMI, 2025CVPR]. It comprises 530 volumetric scans with voxel-wise annotations of maxillofacial anatomy, of which 480 are publicly available for training, while 50, acquired by a different institution, are held out for evaluation via the grand-challenge platform. The dataset provides annotations for 42 classes, including jawbones, left and right inferior alveolar canals, maxillary sinuses, pharynx, upper and lower teeth, bridges, crowns, and implants. For evaluation, we obtained access to the private test set, collected by a different centre on different acquisition machines.
Appendix 0.B Models
We trained a range of model architectures to build a diverse model set for transfer comparison. All of the models trained by us are trained on the training set of a single ‘source’ dataset. We additionally investigated several publicly available models taken from the Toothfairy2 challenge [2024TMI], the Bioimage Model Zoo [ouyang_bioimage_2022] and large generalist models taken from popular repositories (Cellpose-SAM [pachitariu_cellpose-sam_2025], SAM [archit_segment_2025], TotalSegmentator [wasserthal_totalsegmentator_2023]. Information listing the full set of models considered can be found in Sec.˜4.2 in the main paper. In the following section we detail the model training setup.
@@ Appendix 0.B Models
0.B.1 Instance Segmentation Training
For the instance segmentation models trained by us we follow a popular instantiation process [pape_microscopy-based_2021] of predicting cell boundaries and then using seeded watershed and graph partitioning algorithm GASP [bailoni_gasp_2022] to obtain instances.
Appendix 0.C Model Perturbation
@@
Feature Space Perturbations: We relied upon test-time application of spatial DropOut [tompson_efficient_2015] to provide feature space perturbation. We tested several approaches for applying DropOut to our models. Firstly, we tested applying Spatial DropOut only to the bottleneck layer of U-Net style architectures. Our experiments for bottleneck only DropOut are shown on the nuclei datasets both for semantic and instance segmentation (Tabs.˜11(a), 11(b) and 21). Secondly, in addition to the bottle neck layer we also investigated adding DropOut to all layers that connect the encoder to the decoder, namely skip connections in U-Net style networks. Our experiments for bottleneck + skip connection DropOut are shown on the Toothfairy2 dataset for multiclass semantic segmentation (Tab.˜14). Lastly, we investigate applying DropOut uniformly to every layer of the network, this represents the most general and architecture agnostic approach considered. Our experiments for all layer DropOut are shown on the mitochondria semantic segmentation datasets along with the cell instance segmentation datasets (Tabs.˜6(a), 6(b) and 19).
@@
We found that all the considered feature space perturbation approaches were able to perform effectively, as shown in the tables in the following sections. However, it is worth noting that the first approach of applying DropOut to the bottleneck layer only can be sensitive to residual connections. The second approach of applying DropOut to both the bottleneck layer and skip connections works very well on the model sets tested, but does assume that the models have skip connections in the architecture, otherwise it reduces to the first case (bottleneck only) and could suffer the same sensitivity to residual connections. The last approach of applying DropOut to all layers is the most architecture agnostic and resulted in very strong correlation between the CMR metric and performance score.
Appendix 0.D Unsupervised Domain Adaptation (UDA) Approaches
When considering the reuse of pre-trained models users can either directly apply a model with frozen weights, in what we refer to as zero-shot reuse, or after applying some unsupervised domain adaptation (UDA) method to models in order to adapt them to the target data. Application of UDA has the potential to improve model performance on the target data, but requires users to have the resources and know-how to properly make use of existing methods. Furthermore, UDA methods can often be unstable and can even lead to degradation of model performance on the target data [wang_characterizing_2019]. Hence, even if UDA is applied the ranking task still persists, and now the best adapted model must be selected.
In order to investigate ranking of post-UDA models we utilised two common UDA approaches to adapt our pretrained source networks to target data. Firstly, we used a Mean Teacher [tarvainen_mean_2018] approach that relies on a ‘student’ ‘teacher’ paired network setting. The Mean Teacher framework for self supervised training initialises both the student and teacher networks using the pre-trained source model weights and then relies on the ‘teacher’ branch to provide pseudo labels via its predictions. The student branch is trained via gradient descent using a consistency loss to the pseudo labels. Finally, the teacher weights are then updated via an exponential moving average of the student weights. We did not employ any additional pseudo-label selection steps as we found ‘no-selection’ was a sufficient strategy in many cases for successful UDA. In addition, in our experiments we are not actually directly interested in improving the performance of UDA, but rather our real question is: can we rank the final post-UDA models regardless of whether UDA successfully improved target performance or indeed led to model degradation?
The second UDA approach we employed was Adaptive Batch Normalization [li_adaptive_2018] (AdaBN). AdaBN provides a simple and parameter-free approach for UDA and is based on updating the statistics of all Batch Normalization layers in a network to match the target data. This simple approach proved surprisingly effective at improving post-UDA model performance.
For both the Mean Teacher and AdaBN approaches we were able to successfully rank the post-UDA model performance, outcompeting the state-of-the-art baseline Transfer Score [yang_can_2024]. In Tab.˜4, we showed the results for single CMR perturbation strengths, in Tab.˜16(b) we show the CMR correlation to post-UDA F1-score across a range of perturbation strengths.
Appendix 0.E Semantic Segmentation
@@
In the following section we investigate ranking of semantic segmentation models, each of the tables Tabs.˜5(a), 6(a), 7(a), 11(a), 5(b), 6(b), 7(b) and 11(b) shows the correlation scores between our CMR metrics (CMR-EI and CMR-NHD) and the performance F1-score for sets of models applied directly to a single target dataset, across a range of perturbation strengths. In the main paper, due to space constraints, we show the correlation scores for a single perturbation strength averaged over the set of target datasets.
@@
We investigate the dependency of CMR on the strength of perturbation applied for both input and feature space perturbations. The tables show that CMR’s performance is stable across a wide range of perturbations.
0.E.1 Mitochondria Perturbation Sweep
For mitochondria semantic segmentation we investigated ranking a model set with . The models were trained separately using all four Mitochondria datasets (EPFL [lucchi_learning_2013], Hmito [franco-barranco_current_2023], Rmito [franco-barranco_current_2023] and VNC [phelps_reconstruction_2021]) as different sources. The model set comprised of a range of model architectures: 2D U-Net (trained both with and without augmentations), Residual U-Net and UNETR. Hence we built a diverse set of domain-specialist models which we apply to each target dataset in turn and rank the corresponding performance via CMR.
@@ 0.E.1 Mitochondria Perturbation Sweep
To investigate input space perturbations we applied additive Gaussian noise to the inputs between the strengths . We show the results for both CMR-EI (Tab.˜5(a)) and CMR-NHD (Tab.˜5(b)).
@@
For feature space perturbation we investigated applying spatial DropOut equally to all layers of a network, where the strength of the perturbation is controlled by the proportion of feature maps dropped at each layer . We investigated DropOut proportions in the range . We show results for for both CMR-EI (Tab.˜6(a)) and CMR-NHD (Tab.˜6(b)). The tables show that CMR’s performance is stable across a wide range of perturbations.
@@
0.E.2 Nuclei Perturbation Sweep
For nuclei semantic segmentation we investigated ranking a model set with . We built the model set by training a set of specialist 2D U-Net models on seven distinct source datasets: BBBC039 [ljosa_annotated_2012], Go-Nuclear [vijayan_deep_2024], HeLaCytoNuc [de_helacytonuc_2024], Hoechst [arvidsson_annotated_2023], S_BIAD895/SB-895 [von_chamier_democratising_2021], SELMA3D 2024 challenge [leal-taixe_benchmark_2019] and S_BIAD1410 [hawkins_rescu-nets_2024].
We then investigated applying a range of input augmentations; additive Gaussian noise, Gamma correction, Brightness adjustment and Contrast adjustment (as defined in Appendix˜0.C). We investigated additive Gaussian noise with strengths in the range , Gamma correction with strengths in the range , where equals no adjustment, Brightness adjustment with strengths in the range and Contrast adjustment with strengths in the range , where equals no adjustment. We show the results for both CMR-EI (Tabs.˜7(a), 8(a), 9(a) and 10(a)) and CMR-NHD (Tabs.˜7(b), 8(b), 9(b) and 10(b)).
For feature space perturbation we investigated applying spatial DropOut only to the bottle neck layer of networks, where the strength of the perturbation is controlled by the proportion of feature maps dropped at each layer . We investigated DropOut proportions in the range . We show results for for both CMR-EI (Tab.˜11(a)) and CMR-NHD (Tab.˜11(b)). Again over a wide range of perturbation strengths across all considered input and feature space perturbations the correlation scores achieved by our CMR metrics remain largely stable. Although, as discussed in the main paper, Sec.˜3.3, care should be taken to ensure perturbations remain ‘tolerable’ in the extreme perturbation case, too weak perturbation can lead to a loss of correlation between consistency based metrics and transfer performance. For example, the first row of Tab.˜7(a) for the DSB2018 target dataset shows the weakest level of Gaussian input noise () can cause the CMR-NHD metric to become unstable on DSB2018.
| Metric | BBBC039 | DSB2018 | Hoechst | SB-895 | |||||
@@ 0.E.2 Nuclei Perturbation Sweep
| Metric | BBBC039 | DSB2018 | Hoechst | SB-895 | |||||
@@
| Metric | BBBC039 | DSB2018 | Hoechst | SB-895 | |||||
@@
| CMR-EI Brt, | K | 0.71 | (*) | 0.43 | (0.25) | 0.52 | (0.16) | 0.60 | (0.12) |
| S | 0.86 | (*) | 0.64 | (0.15) | 0.71 | (0.08) | 0.77 | (0.08) | |
| P | 0.99 | (**) | 0.91 | (**) | 0.94 | (**) | 0.96 | (**) | |
| CMR-EI Brt, | K | 0.81 | (*) | 0.62 | (0.08) | 0.81 | (*) | 0.73 | (0.07) |
| S | 0.93 | (*) | 0.75 | (0.07) | 0.93 | (**) | 0.83 | (0.05) | |
| P | 0.99 | (**) | 0.99 | (**) | 0.94 | (**) | 0.99 | (**) | |
| CMR-EI Brt, | K | 0.71 | (*) | 0.62 | (0.06) | 0.81 | (*) | 0.73 | (0.06) |
| S | 0.86 | (*) | 0.75 | (0.07) | 0.93 | (*) | 0.83 | (0.07) | |
| P | 0.99 | (**) | 0.99 | (**) | 0.94 | (**) | 0.99 | (**) | |
| Metric | BBBC039 | DSB2018 | Hoechst | SB-895 | |||||
| pval. | pval. | pval. | pval. | ||||||
| CMR-NHD Brt, | K | 0.52 | (0.12) | 0.71 | (*) | 0.71 | (*) | 0.47 | (0.25) |
| S | 0.71 | (0.08) | 0.79 | (0.05) | 0.86 | (*) | 0.54 | (0.29) | |
| P | 1.00 | (**) | 0.82 | (*) | 0.92 | (**) | 0.82 | (0.05) | |
| CMR-NHD Brt, | K | 0.71 | (*) | 0.62 | (0.07) | 0.71 | (*) | 0.73 | (0.05) |
| S | 0.86 | (*) | 0.71 | (0.08) | 0.86 | (*) | 0.83 | (0.06) | |
| P | 1.00 | (**) | 0.98 | (**) | 0.90 | (*) | 0.99 | (**) | |
| CMR-NHD Brt, | K | 0.71 | (*) | 0.71 | (*) | 0.90 | (*) | 0.73 | (0.07) |
| S | 0.86 | (*) | 0.82 | (*) | 0.96 | (**) | 0.83 | (0.07) | |
| P | 1.00 | (**) | 0.99 | (**) | 0.93 | (**) | 0.99 | (**) | |
| Metric | BBBC039 | DSB2018 | Hoechst | SB-895 | |||||
| pval. | pval. | pval. | pval. | ||||||
| CMR-EI Ctr, | K | 0.81 | (*) | 0.33 | (0.35) | 0.52 | (0.16) | 0.47 | (0.32) |
| S | 0.93 | (*) | 0.50 | (0.26) | 0.68 | (0.12) | 0.71 | (0.11) | |
| P | 0.95 | (**) | 0.91 | (**) | 0.94 | (**) | 0.98 | (**) | |
| CMR-EI Ctr, | K | 0.71 | (*) | 0.33 | (0.36) | 0.52 | (0.15) | 0.47 | (0.26) |
| S | 0.86 | (*) | 0.50 | (0.27) | 0.68 | (0.12) | 0.71 | (0.12) | |
| P | 0.98 | (**) | 0.95 | (**) | 0.95 | (**) | 0.98 | (**) | |
| CMR-EI Ctr, | K | 0.62 | (0.09) | 0.43 | (0.28) | 0.62 | (0.08) | 0.47 | (0.27) |
| S | 0.71 | (0.10) | 0.64 | (0.13) | 0.75 | (0.09) | 0.71 | (0.13) | |
| P | 0.98 | (**) | 0.97 | (**) | 0.94 | (**) | 0.98 | (**) | |
| CMR-EI Ctr, | K | 0.62 | (0.07) | 0.43 | (0.26) | 0.62 | (0.09) | 0.47 | (0.29) |
| S | 0.71 | (0.11) | 0.64 | (0.15) | 0.75 | (0.05) | 0.71 | (0.15) | |
| P | 0.99 | (**) | 0.97 | (**) | 0.95 | (**) | 0.98 | (**) | |
| Metric | BBBC039 | DSB2018 | Hoechst | SB-895 | |||||
| pval. | pval. | pval. | pval. | ||||||
| CMR-NHD Ctr, | K | 0.81 | (*) | 0.43 | (0.23) | 0.52 | (0.12) | 0.73 | (0.05) |
| S | 0.93 | (*) | 0.54 | (0.20) | 0.71 | (0.09) | 0.83 | (0.05) | |
| P | 0.94 | (**) | 0.91 | (**) | 0.92 | (**) | 0.99 | (**) | |
| CMR-NHD Ctr, | K | 0.81 | (*) | 0.43 | (0.25) | 0.81 | (*) | 0.60 | (0.12) |
| S | 0.93 | (*) | 0.54 | (0.24) | 0.93 | (*) | 0.77 | (0.09) | |
| P | 0.98 | (**) | 0.96 | (**) | 0.97 | (**) | 0.98 | (**) | |
| CMR-NHD Ctr, | K | 0.81 | (*) | 0.43 | (0.27) | 0.71 | (*) | 0.73 | (0.07) |
| S | 0.93 | (**) | 0.54 | (0.21) | 0.89 | (*) | 0.83 | (0.07) | |
| P | 0.99 | (**) | 0.97 | (**) | 0.92 | (**) | 0.99 | (**) | |
| CMR-NHD Ctr, | K | 0.81 | (*) | 0.43 | (0.24) | 0.71 | (0.05) | 0.60 | (0.14) |
| S | 0.93 | (*) | 0.64 | (0.12) | 0.89 | (*) | 0.77 | (0.08) | |
| P | 0.99 | (**) | 0.99 | (**) | 0.91 | (**) | 0.99 | (**) | |
| Metric | BBBC039 | DSB2018 | Hoechst | SB-895 | |||||
| pval. | pval. | pval. | pval. | ||||||
| CMR-EI Gamma, | K | 0.71 | (*) | 0.52 | (0.12) | 0.81 | (**) | 0.60 | (0.12) |
| S | 0.86 | (*) | 0.68 | (0.13) | 0.89 | (*) | 0.77 | (0.09) | |
| P | 0.98 | (**) | 0.96 | (**) | 0.98 | (**) | 0.97 | (**) | |
| CMR-EI Gamma, | K | 0.71 | (*) | 0.52 | (0.14) | 0.71 | (*) | 0.47 | (0.26) |
| S | 0.86 | (*) | 0.68 | (0.13) | 0.82 | (*) | 0.71 | (0.16) | |
| P | 0.99 | (**) | 0.96 | (**) | 0.98 | (**) | 0.96 | (**) | |
| CMR-EI Gamma, | K | 0.71 | (*) | 0.52 | (0.13) | 0.62 | (0.06) | 0.47 | (0.26) |
| S | 0.86 | (*) | 0.68 | (0.12) | 0.75 | (0.08) | 0.71 | (0.14) | |
| P | 0.98 | (**) | 0.94 | (**) | 0.97 | (**) | 0.97 | (**) | |
| CMR-EI Gamma, | K | 0.81 | (*) | 0.52 | (0.16) | 0.81 | (*) | 0.73 | (0.06) |
| S | 0.93 | (**) | 0.68 | (0.14) | 0.89 | (*) | 0.83 | (0.07) | |
| P | 0.96 | (**) | 0.94 | (**) | 0.98 | (**) | 0.96 | (**) | |
| Metric | BBBC039 | DSB2018 | Hoechst | SB-895 | |||||
| pval. | pval. | pval. | pval. | ||||||
| CMR-NHD Gamma, | K | 0.71 | (*) | 0.62 | (0.05) | 0.90 | (*) | 0.87 | (*) |
| S | 0.86 | (*) | 0.75 | (0.06) | 0.96 | (**) | 0.94 | (*) | |
| P | 0.97 | (**) | 0.91 | (**) | 0.99 | (**) | 0.95 | (**) | |
| CMR-NHD Gamma, | K | 0.62 | (0.07) | 0.52 | (0.14) | 0.81 | (*) | 0.87 | (*) |
| S | 0.75 | (0.06) | 0.68 | (0.13) | 0.93 | (*) | 0.94 | (*) | |
| P | 0.97 | (**) | 0.92 | (**) | 1.00 | (**) | 0.95 | (**) | |
| CMR-NHD Gamma, | K | 0.62 | (0.08) | 0.43 | (0.26) | 1.00 | (**) | 0.73 | (0.05) |
| S | 0.82 | (0.05) | 0.57 | (0.19) | 1.00 | (**) | 0.83 | (0.08) | |
| P | 0.93 | (**) | 0.88 | (*) | 1.00 | (**) | 0.83 | (*) | |
| CMR-NHD Gamma, | K | 0.71 | (*) | 0.62 | (0.06) | 0.81 | (*) | 1.00 | (**) |
| S | 0.79 | (*) | 0.75 | (0.06) | 0.93 | (*) | 1.00 | (*) | |
| P | 0.84 | (*) | 0.89 | (*) | 1.00 | (**) | 0.80 | (0.06) | |
| Metric | BBBC039 | DSB2018 | Hoechst | SB-895 | |||||
| pval. | pval. | pval. | pval. | ||||||
| CMR-EI DropOut 0.05 | K | 0.71 | (*) | 0.43 | (0.24) | 0.52 | (0.18) | 0.60 | (0.13) |
@@
| Metric | BBBC039 | DSB2018 | Hoechst | SB-895 | |||||
@@
| CMR-NHD DropOut 0.05 | K | 0.71 | (*) | 0.43 | (0.23) | 0.52 | (0.15) | 0.87 | (*) |
@@
0.E.3 Baseline Transferability Metrics
We also show the correlation scores per dataset for all the baseline transferability metrics. Tab.˜12 shows the results for the four mitochondria datasets (EPFL [lucchi_learning_2013], Hmito [franco-barranco_current_2023], Rmito [franco-barranco_current_2023] and VNC [phelps_reconstruction_2021]) and Tab.˜13 shows the results for the four nuclei datasets (BBBC039 [ljosa_annotated_2012], DSB2018 [caicedo_nucleus_2019], Hoechst [arvidsson_annotated_2023], S_BIAD895 [von_chamier_democratising_2021]). In the main paper due to space constraints we showed the correlation scores averaged within the dataset groups.